OpenStudy (anonymous):

Write the expression in the form A log(x)+B log(x+1)+C log(x-2) using the laws of logarithms

OpenStudy (anonymous):

\[\large \log((x^{3}\sqrt{x+1})/(x^2-4x+4))\]

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

Log ( A/ B ) = log A - Log B

OpenStudy (anonymous):

i get this so far \[\large \log(x^{3})+\log(\sqrt{x+1})-\log(x^{2}-4x+4)\]

OpenStudy (anonymous):

dunno if its right

OpenStudy (anonymous):

Bring the power down: 3 log x + (1/2) log ( x +1) - 2log ( x -2)

OpenStudy (anonymous):

Hmm, so you factored x^2-4x+4 as (x-2)^2 and then put the 2 in front?

OpenStudy (anonymous):

Yes, always attempt to convert into simplest form :)

OpenStudy (anonymous):

I see, that was way simpler than i thought. Thanks

OpenStudy (anonymous):

:)